OpenStudy (anonymous):

A sports car starting from rest accelerates uniformly at 2m/s^2,what is its speed after 10s?how far has it traveled after 15s?and after how long will the car be 500m from its starting point? show the sulotion

OpenStudy (anonymous):

v-vi=at v-0=(2m/s^2)10=20m/s^2 after t=15sec: s=1/2 vt s=1/2 (20m/s)15s=150m and using s=1/2 vt 500m= 1/2 (20)t 500/10=t=50 sec

OpenStudy (anonymous):

v-vi=at v-0=(2m/s^2)10=20m/s^2 after t=15sec: s=1/2 vt s=1/2 (20m/s)15s=150m and using s=1/2 vt 500m= 1/2 (20)t 500/10=t=50 sec